OpenStudy (cheddarskittles):

Show work? 4b+8 _______=10 -2

OpenStudy (anonymous):

-2b-4=10 -2b=14 b=-7 Happy birthday

OpenStudy (cheddarskittles):

Thanks. I have about 7-8 more questions can you help?

OpenStudy (cheddarskittles):

@Robert136

OpenStudy (anonymous):

Post them new

OpenStudy (cheddarskittles):

ok